# Harnessing The Collective Wisdom: Fusion Learning Using Decision Sequences From Diverse Sources

### v2

@@ 5 Real data illustration
Figure 4: Left: frequency distribution of $n_{i}$’s. Right: heatmap of the $\log$ evidence indices for $100$ randomly sampled genes across the $d$ studies. White indicates genes not analyzed, while shades of brown represent evidence indices of 00, indicating failure to reject.
The sixth column of Table 3 reports the number of rejections for each of these studies and the last column presents the evidence against each rejected null hypothesis in study $j$. It is interesting to see that study 5 (Varambally et al., 2005) receives the highest evidence for its rejected hypotheses, which is not surprising given the large weight $w_{5}$ that each of its relatively small number of rejections receives. In contrast, study 8 (Wallace et al., 2008) has the smallest non-zero evidence which is driven by the largest number of rejections reported in this study. The right panel of Figure 4 presents a heatmap of the $\log$ evidence indices for $100$ randomly sampled genes across the $d$ studies. Here the white shade represents a gene not analyzed by the study while the shade of brown represents an evidence index of 00 which corresponds to a failure to reject the underlying null hypothesis. The heterogeneity across the $d$ studies is evident through the different magnitudes of the evidence indices constructed for each study.
Table 4: Distribution of rejection overlaps across $7$ studies.
@@
5.1 Ranking and thresholding using IRT
In this section we aggregate the evidence indices using Equation (3) and threshold the ordered aggregated evidences using the $e-$BH procedure at $\alpha=0.1$. We recall that this thresholding scheme guarantees valid FDR control under unknown and arbitrary dependence between the aggregated evidences.
@@ 5.1 Ranking and thresholding using IRT
The top-left panel of Figure 5 presents a histogram of the log-transformed non-zero aggregated evidence from IRT, while the top-right panel plots the top 25 genes with respect to their aggregated evidence, colored and shape-coded by the number of times the corresponding gene was analyzed across the $d$ studies. Interestingly, the top second and third genes have $n_{i}=2$ and $3$, respectively, suggesting that apart from the number of times a particular null hypothesis is analyzed across the $d$ studies, the magnitude of the study-specific evidence indices also play a key role in the overall ranking. To put this into perspective, the bottom-right panel of Figure 5 presents the top 25 genes with respect to their aggregated evidence from the P2E framework introduced in Section 4.3. In stark contrast to IRT, here the top 25 genes have $n_{i}\geq 7$. Furthermore, both the left and right panels of Figure 5 suggest that $e_{i}^{\texttt{P2E}}$ can be substantially larger in magnitude than $e_{i}^{\tt agg}$ particularly when one of the studies rejects the null hypothesis with an astronomically small $p-$value.
@@
5.2 Ranking and thresholding using IRT
Here we aggregate the evidence indices using the scheme discussed in Section 3.2 and threshold the ordered aggregated evidences using the $e-$BH procedure at $\alpha=0.005$. We note that this thresholding scheme guarantees valid FDR control under exchangeability of the study-specific summary statistics (Definition 2), symmetry of the study-specific decision rule (Definition 3) and independence of the $n_{i}$ summary statistics for each testing problem.
@@ 5.2 Ranking and thresholding using IRT
6 Numerical experiments
In this section we assess the empirical performances of IRT, IRT and IRT H on simulated data. We consider eight simulation scenarios with $m=1000$ and test $H_{0i}:\mu_{i}=0~{}vs~{}H_{1i}:\mu_{i}\neq 0$, where
$\mu_{i}\stackrel{{\scriptstyle\rm i.i.d.}}{{\sim}}0.8\cdot\delta_{\{0\}}+0.1\cdot\mathcal{N}(3,1)+0.1\cdot\mathcal{N}(-3,1),$
and $\delta_{\{a\}}$ denotes a point mass at $a$. In each scenario, study $j$ uses data $X_{ij}$, to be specified subsequently, to conduct $m_{j}$ tests and reports the corresponding decisions $\bm{\delta}_{j}$ obtained from the BH procedure with control level $\alpha_{j}$. The empirical performances of IRT and its derivatives are compared against three alternative procedures: (i) Fisher, which pools the study specific $p-$values using Fisher’s method (Fisher, 1948), and then applies the BH procedure on the pooled $p-$value sequence for FDR control, (ii) the P2E procedure from Section 4.3 and (iii) the P2E procedure, also from Section 4.3. For IRT H, we use the following scheme across all our simulation settings: the IRT aggregation scheme (Equation (3)) is employed for the first $d_{1}=\lfloor d/2\rfloor$ studies and the IRT aggregation scheme (Equation (7)) is used for the remaining $d_{2}=d-d_{1}$ studies. When the null distribution of the test statistic is correctly specified and the corresponding $p-$values are independent, we expect Fisher to exhibit higher power than $e-$value procedures, such as IRT and P2E. Nevertheless, in such settings Fisher provides a practical benchmark for assessing the empirical performances of P2E, P2E, IRT, IRT and IRT H, where the last three procedures rely only the binary decision sequences $\bm{\delta}_{j}$.
Figure 7: FDP and ETP comparisons for Scenario 1.
Scenario 1 (Independent studies). We begin by considering $d$ independent studies. Specifically, we let $X_{ij}\mid\mu_{i},\sigma_{j}\stackrel{{\scriptstyle\rm ind.}}{{\sim}}\mathcal{N}(\mu_{i},\sigma_{j}^{2}),~{}\sigma_{j}\stackrel{{\scriptstyle\rm i.i.d.}}{{\sim}}{\rm Unif}(0.75,2),m_{j}=m$, $\alpha_{j}=0.01$, $\alpha=0.005$ and vary $d$ from $5$ to $10$. Figure 7 presents the average FDP and the ETP across 200 Monte Carlo repetitions. We make several observations. First, while all methods control the FDR at $\alpha$, Fisher, unsurprisingly, has the highest power across all values of $d$ and is followed by P2E. Second, IRT H is more powerful than IRT. In fact, the latter exhibits no power since $\alpha<\alpha_{j}$ for all $j\in[d]$ in this setting, further reinforcing the discussion in Section 3.1 and Remark 1. However, this is not the case when the studies are correlated and $\alpha>\max_{j\in[d]}\alpha_{j}$, as scenarios 3 and 4 demonstrate. Third, IRT is less powerful than P2E. The explanation of this phenomenon lies in the observation that in this setting P2E employs optimal $p-$values as the underlying null distribution of the test statistic is fully known. However, when the null distribution is misspecified P2E and P2E may not even provide valid FDR control. Scenarios 7 and 8, reported later in this section, consider such a setting. Finally, IRT is more powerful than IRT H. This is expected since in this setting the conditions of Theorem 3 hold and the aggregation scheme of Equation (7) results in a more powerful procedure than IRT H.
Figure 8: FDP and ETP comparison for Scenario 2.
Scenario 2 (Correlated studies I). The data are generated according to Scenario 1 with $\sigma_{j}=1,d=10,\alpha_{j}=0.01,\alpha=0.005$ but we introduce correlation across the first $d_{1}=\lfloor d/2\rfloor$ studies. In particular, we generate $(X_{i1},\ldots,X_{i\lfloor d/2\rfloor})$ from a multivariate normal distribution and set $\text{Corr}(X_{ij},X_{ik})=\rho$ for all $(j,k)\in\{1,\ldots,\lfloor d/2\rfloor\}$, $j\neq k$, where $\rho\in\{-0.1,0,0.1,0.3,0.5,0.7,0.9\}$. Figure 8 reports the average FDP and the ETP for various methods. In this scenario, the first $d_{1}$ test statistics and the corresponding $p-$values for each hypothesis are not independent unless $\rho=0$, thus violating the conditions of Theorem 3. Consequently, P2E, IRT and Fisher no longer guarantee FDR control. Indeed, the left panel of Figure 8 reveals that P2E and Fisher fail to control the FDR at $0.5\%$ for large $\rho$ and therefore do not appear in the plot for some values of $\rho$. While IRT appears to control the FDR, it has no theoretical support for FDR guarantee in this setting. From the right panel, we find that IRT H is the most powerful procedure in this setting that also provably controls FDR.
Figure 9: FDP and ETP comparison for Scenario 3.
Scenario 3 (Correlated studies II). We continue to borrow the setting from Scenario 1 with $\sigma_{j}=1,~{}d=10,~{}\alpha_{j}=0.01,~{}\alpha=0.05$ and introduce correlation across all $d$ studies. Specifically, we let $\text{Corr}(X_{ij},X_{ik})=\rho$ for all $j\neq k$, where $\rho\in\{-0.1,0,0.1,0.3,0.5,0.7,0.9\}$. Figure 9 reports the average FDP and the ETP for various methods. In this scenario, the $d$ test statistics and the corresponding $p-$values for each hypothesis are not independent unless $\rho=0$. Thus, P2E, IRT, IRT H and Fisher have no theoretical support for FDR control at level $\alpha$. As observed in Scenario 2, P2E and Fisher continue to fail to control the FDR at $5\%$ for large $\rho$ and therefore do not appear in the left panel of Figure 9 for some values of $\rho$. In fact IRT is the most powerful procedure in this setting that also provably controls FDR. While several methods for pooling exchangeable $p-$values are available (see for instance Vovk and Wang (2020)), all of them rely on some prior knowledge regarding the strength of the dependence between the $p-$values, which may not be available in practice.
Scenario 4 (Correlated studies and dependent test statistics). We generate the data from Scenario 1 with $\sigma_{j}=1,\alpha_{j}=0.01,\alpha=0.02$ and introduce correlation across the studies as well as the test statistics. In particular, we let $\text{Corr}(X_{ij},X_{ik})=0.7,~{}j\neq k$ and $\text{Corr}(X_{ij},X_{rj})=0.5,i\neq r$. In this setting too P2E, IRT, IRT H and Fisher no longer enjoy theoretical guarantees for FDR control. Figure 10 reports the average FDP and the ETP for various methods as $d$ varies from $5$ to $10$. We see a similar pattern as Figure 9 where P2E and Fisher fail to control the FDR at $2\%$ while IRT controls the FDR and exhibits higher power than P2E.
Figure 10: FDP and ETP comparison for Scenario 4.
Figure 11: FDP and ETP comparison for Scenario 5.
Scenario 5 (Varying $m_{j}$ and $\alpha_{j}$). In this scenario we go back to the setting of independent studies. The data are generated according to Scenario 1 with $\sigma_{j}=1,m=1000,d=10$ and $\alpha=0.03$, but we vary $(m_{j},\alpha_{j})$ for the $d$ studies. To vary $m_{j}$, we set $m_{(1)}=\max\{m_{1},\ldots,m_{d}\}=900$ and consider the ratio $\eta=\min\{m_{1},\ldots,m_{d}\}/m_{(1)}$. For a given choice of $\eta$, we first sample $m_{1},\ldots,m_{d}$ uniformly from $[\lceil m_{(1)}\eta\rceil,m_{(1)}]$ with replacement and then for each $j$, $m_{j}$ hypotheses are chosen at random from the $m$ hypotheses without replacement. We set $\alpha_{j}\in\{0.05,0.03,0.01\}$ according to $m_{j}\leq 600$, $m_{j}\in(600,800]$ or $m_{j}>800$, respectively. Thus, in this setting studies with a higher $m_{j}$ receive a larger weight on their rejections. Figure 11 reports the average FDP and the ETP for various methods as $\eta$ varies over $[0.1,0.5]$. We find that both IRT H and IRT exhibit higher power as $\eta$ increases and dominate IRT in power for all values of $\eta$. Furthermore, when $\eta$ is large, $m_{j}$ is large and studies receive a relatively higher weight $w_{j}$ on their rejections, which leads to an improved power in this setting.
Figure 12: FDP and ETP comparison for Scenario 6.
Scenario 6 (Sensitivity to $\alpha$). In this scenario we report the performance of the six procedures for varying $\alpha$. The data are generated according to Scenario 1 with $\sigma_{j}=1,m=1000,d=10$, but with $(m_{j},\alpha_{j})$ varying across the $d$ studies. To vary $m_{j}$, we first sample $m_{1},\ldots,m_{d}$ uniformly from $\{2000,2500,3000\}$ with replacement and then set $\alpha_{j}\in\{0.05,0.03,0.01\}$ according to $m_{j}=2000$, $2500$ or $3000$, respectively. Thus, in this setting $m<m_{j}$ and studies with a higher $m_{j}$ continue to receive a larger weight on their rejections. Finally, we uniformly sample the $m$ hypotheses without replacement from $\{1000,\ldots,3000\}$ which ensures that $n_{i}\geq 4$ for all $i\in\{1,\ldots,m\}$. Figure 12 reports the average FDP and the ETP for various methods as $\alpha$ varies over $\{0.001,0.003,0.005,0.01,0.03,0.05,0.1\}$. Several observations are in order. First, the left panel reveals that all methods control FDR at level $\alpha$. Second, the right panel reveals that IRT and IRT H are powerful even when $\alpha<\min_{j\in[d]}\alpha_{j}$ while IRT exhibits power only for $\alpha\geq 0.05$, thus exemplifying the discussion in Section 3.1. Finally, when $\alpha\geq 0.05$ IRT is more powerful than P2E as observed in scenarios 3 and 4.
Figure 13: FDP and ETP comparisons for Scenario 7.
Scenario 7 (misspecified null distribution: I). We consider two settings where the null distribution of the test statistic is misspecified and the corresponding $p-$values are no longer optimal. For setting 1 we let $m_{j}=m,\alpha_{j}=0.05,\alpha=0.01$ and $d=10$. For agent $j$, the summary statistics $X_{ij}=\mu_{i}+Y_{ij}$ where $Y_{ij}$ are i.i.d. central $t-$distributed random variables with $\nu>2$ degrees of freedom. To test $H_{0i}:\mu_{i}=0~{}vs~{}H_{1i}:\mu_{i}\neq 0$, the $p-$values are calculated assuming a standard Normal distribution for the scaled summary statistics, i.e, $p_{ij}=2\Phi(-|X_{ij}|/\sigma)$ where $\sigma=\sqrt{\nu/(\nu-2)}$. Finally, all procedures rely on these $p-$values for inference. Figure 13 reports the average FDP and the ETP for various methods as $\nu$ varies over $\{5,10,15,30,50,100\}$. We find that while Fisher, P2E and P2E fail to control the FDR at $\alpha$ for small $\nu$, IRT and IRT H control the FDR for all values of $\nu$.
Setting 2 borrows the design from Setting 1 but allows the studies to be correlated, i.e, $\text{Corr}(X_{ij},X_{ik})=0.7$ for all $j\neq k$, and sets $\alpha_{j}=0.01,\alpha=0.02$. Since the conditions of Theorem 3 do not hold in this setting and the corresponding $p-$values are not independent, we exclude Fisher, P2E, IRT and IRT H from our comparisons. Figure 13 reports the results of this setting and reveals that P2E does not control the FDR at level $\alpha$ for $\nu=5$. Furthermore, IRT is far more powerful than P2E in this setting for all but the smallest value of $\nu$.
Figure 14: FDP and ETP comparisons for Scenario 8.
Scenario 8 (misspecified null distribution: II). Here we consider two additional settings where the null distribution is misspecified. For setting 1 $m_{j}=m,\alpha_{j}=0.05,\alpha=0.01$ and $d=10$. For agent $j$, the summary statistics $X_{ij}=\mu_{i}+Y_{ij}$ where $Y_{ij}\stackrel{{\scriptstyle i.i.d}}{{\sim}}(1-q)t_{5}+qN(0,1)$ and $q\in[0,1]$. Here $t_{5}$ is a central $t-$distributed random variable with $5$ degrees of freedom. The corresponding $p-$values are calculated assuming a standard Normal distribution for the scaled summary statistics, i.e, $p_{ij}=2\Phi(-|X_{ij}|/\sigma)$ where $\sigma=\sqrt{q+(1-q)5/3}$. Figure 14 reports the average FDP and the ETP for various methods as $q$ varies over $\{0.1,0.3,0.5,0.7,0.9\}$. The three $p-$value based procedures fail to control the FDR at $\alpha$ for all values of $q$. In contrast, IRT and IRT H control the FDR and are also powerful. Setting 2 borrows the design from Setting 1 but allows the studies to be correlated, i.e, $\text{Corr}(X_{ij},X_{ik})=0.7$ for all $j\neq k$, and sets $\alpha_{j}=0.01,\alpha=0.02$. We continue to exclude Fisher, P2E, IRT H and IRT from our comparisons in this setting. Figure 14 reports the results of this setting and reveals that IRT is a relatively better procedure than P2E, both in terms of FDR control and power.
Overall, the results of scenarios 7 and 8 reveal that when the $p-$values are misspecified, procedures that rely directly on the magnitude of the $p-$values for pooling inferences may fail to control FDR at the desired level. In contrast, IRT, IRT H and IRT are relatively robust to such misspecification. In fact in scenarios 7 and 8 they not only control FDR across all levels of misspecification but also retain power. We, however, caution that IRT and its derivatives are not without limitations. A fundamental requirement underlying the proposed procedures is that the studies control FDR at their desired levels. If the study level FDR control is severely violated, either due to misspecified $p-$values or a faulty multiple testing methodology, then our framework may fail to control the overall FDR. Nevertheless, even in those settings the proposed procedures will be relatively less liberal in making false discoveries than methods that rely directly on the magnitudes of misspecified $p-$values for pooling inferences.
7 IRT for alternative forms of Type I error control
Sections 2 and 3 present IRT for integrative FDR control. Here, we extend IRT to alternative forms of Type I error control.
7.1 IRT for $k$-Family-wise error rate ($k$-FWER) control
$\mathbb{P}\left(\sum_{i\in\mathcal{H}_{0j}}\delta_{ij}\geq k_{j}\right)\leq\alpha_{j},\quad\forall j\in[d].$ (17)
  1. 1.
    If study $j$ satisfies Equation (17) then the evidence indices in Equation (18) are generalized $e-$values. To see this, let $V_{j}$ denote the number of false rejections made by study $j$. Then,
    We note that if additional information, such as what procedure study $j$ used to control $k_{j}$-FWER, is available then it becomes feasible to devise more powerful generalized $e-$values. For example, if Bonferroni procedure is used (i.e. $H_{0i}$ is rejected by study $j$ if and only if its $p-$value is $\leq k_{j}\alpha_{j}/m_{j}$), then we can verify that
    $e_{ij}=\dfrac{m_{j}\delta_{ij}}{\alpha_{j}k_{j}},\quad\forall i\in\mathcal{M}_{j},$
    is also a generalized $e-$value. To see this, observe that
    $\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0j}}e_{ij}\right]\leq\frac{m_{j}}{k_{j}\alpha_{j}}\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0j}}\delta_{ij}\right]=\dfrac{m_{j}}{k_{j}\alpha_{j}}\sum_{i\in\mathcal{H}_{0j}}\mathbb{P}(\delta_{ij}=1)\leq\dfrac{m_{j}}{k_{j}\alpha_{j}}\sum_{i\in\mathcal{H}_{0j}}\dfrac{k_{j}\alpha_{j}}{m_{j}}\leq m_{j}.$
  2. 2.
    Evidence aggregation. Since the evidence indices in Equation (18) are generalized $e-$values, Equation (3) provides the evidence aggregation scheme in this setting and Theorem 2 guarantees that these aggregated evidences continue to be generalized $e-$values associated with $\mathcal{H}_{0}$. Furthermore, if the conditions of Theorem 3 hold then Equation (7) represents the aggregated evidence indices.
  3. 3.
    $k$-FWER control. Given the generalized $e-$values $\mathbf{e}^{\tt agg}=\{e_{1}^{\tt agg},\ldots,e_{m}^{\tt agg}\}$ from Step 2 above, IRT rejects $H_{0i}$ if and only if $e_{i}^{\tt agg}\geq m/(\alpha k)$. This procedure controls the $k$-FWER at level $\alpha$ since
7.2 IRT for Per-family error rate (PFER) control
  1. 1.
    Evidence construction. We consider the evidence index
    $e_{ij}=\dfrac{m_{j}\delta_{ij}}{k_{j}},\quad\forall i\in\mathcal{M}_{j}.$ (19)
    It is then straightforward to check that $\mathbf{e}_{j}=\{e_{ij}\}_{i\in\mathcal{M}_{j}}$ are a set of generalized $e-$value associated with $\mathcal{H}_{0j}$.
  2. 2.
    Evidence aggregation. The evidence indices in Equation (19) are generalized $e-$values. Therefore, Equation (3) continues to provide the evidence aggregation scheme in this setting and Theorem 2 guarantees that these aggregated evidences are generalized $e-$values associated with $\mathcal{H}_{0}$. Furthermore, if the conditions of Theorem 3 hold then Equation (7) represents the aggregated evidence indices.
  3. 3.
    $k$-PFER control. Given the Genralized $e-$values $\mathbf{e}^{\tt agg}=\{e_{1}^{\tt agg},\ldots,e_{m}^{\tt agg}\}$ from Step 2 above, IRT rejects $H_{0i}$ if and only if $e_{i}^{\tt agg}\geq m/k$. This procedure controls the PFER at level $k$ since
    $\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}}\mathbb{I}\Big{(}e_{i}\geq m/k\Big{)}\right]\leq\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}}\mathbb{I}\Big{(}e_{i}k/m\geq 1\Big{)}\right]\leq\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}}(e_{i}k/m)\right]\leq\frac{k}{m}\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}}e_{i}\right]\leq k.$
Remark 4.
The IRT framework can be used for integrative inference even when studies employ different type I error control metrics. Suppose, for instance, that $d_{1}$ studies control $k-$FWER, $d_{2}$ control PFER and the remaining $d_{3}$ studies control FDR at desired levels, where $d_{i}\subset[d],i=1,2,3$, $\cap_{i=1}^{3}d_{i}=\emptyset$ and $\cup_{i=1}^{3}d_{i}=[d]$. Denote $e_{ij}^{d_{1}},j\in d_{1},$ as the evidence indices for the $d_{1}$ studies from Equation (18). Similarly, $e_{ij}^{d_{2}}$ and $e_{ij}^{d_{3}}$ denote, respectively, the evidence indices for studies $d_{2}$ and $d_{3}$ from equations (19) and (1). Then $e_{i}=m\left\{\sum_{r=1}^{3}(\sum_{j\in d_{r}}e_{ij}^{d_{r}}/\sum_{j\in d_{r}}m_{j})\right\},~{}i\in\mathcal{M}$, are a set of generalized $e-$values under $\mathcal{H}_{0}$, and the $e-$BH procedure can be applied to $\{e_{i}\}_{i\in\mathcal{M}}$ if, for example, FDR control is the goal. Furthermore, both IRT and IRT H are also applicable in this setting if prior knowledge regarding data-sharing or the use of auxiliary side information for multiple testing is available for the $d$ studies.
8 Conclusion
IRT is a general framework for fusion learning in multiple testing that operates on the binary decision sequences available from diverse studies and conducts integrative inference on the common parameter of interest. It guarantees an overall FDR control under arbitrary dependence between the aggregated evidence indices as long as the studies control their FDR at the desired levels. Under additional assumptions on the data generating process for each study (see Theorem 3), IRT and IRT H provide more powerful procedures for pooling inferences than IRT, while controlling the FDR at the same level. Furthermore, the numerical experiments in Section 6 suggest that when the $p-$values are misspecified, procedures that rely directly on the magnitude of these $p-$values for pooling inferences may fail to control FDR at the desired level. In contrast, IRT, IRT H and IRT are relatively robust to such misspecification and also retain power. The proposed framework can be extended to other popular forms of Type I error control, such as FWER control, as demonstrated in Section 7. We conclude this article with a discussion on several avenues for further investigation.
[…]
Accumulation bias in meta-analysis - In the context of meta-analysis for $m=1$ hypothesis, Schure and Grünwald (2019) recently demonstrate that due to the accumulating nature of scientific evidence dependencies arise between study size $d$ or meta-analysis timing and the results of the meta-analysis. This dependence introduces bias in the estimates, termed accumulation bias, and invalidates the sampling distribution of the test statistic, which is often assumed to be standard Normal. Consequently, Type I errors are inflated as the underlying $p-$values are sub-optimal. The authors recommend using a framework involving likelihood ratios for constructing valid tests as no valid $p-$valued based procedure is feasible in this setting. Note that likelihood ratios are $e-$values and so are the aggregated evidence indices underlying IRT. Therefore, an interesting extension of our framework will be to evaluate the performance of IRT for meta-analysis of $m\geq 1$ hypotheses in the presence of accumulation bias.
9 Acknowledgments
We thank Professor Aaditya Ramdas for his valuable insights on an earlier version of the manuscript. T. Banerjee’s research was supported in part by the University of Kansas General Research Fund #2003743. B. Gang’s research was supported by the National Natural Science Foundation of China grant 12201123. All errors are our own.

### v3

@@ Appendix G Real data illustration
Figure 16: Left: frequency distribution of $n_{i}$’s. Right: heatmap of the $\log$ evidence indices for $100$ randomly sampled genes across the $d$ studies. White indicates genes not analyzed, while shades of brown represent evidence indices of 0, indicating failure to reject.
The sixth column of Table 2 reports the number of rejections for each of these studies and the last column presents the evidence against each rejected null hypothesis in study $j$. It is interesting to see that study 5 (Varambally et al., 2005) receives the highest evidence for its rejected hypotheses, which is not surprising given the large weight $w_{5}$ that each of its relatively small number of rejections receives. In contrast, study 8 (Wallace et al., 2008) has the smallest non-zero evidence which is driven by the largest number of rejections reported in this study. The right panel of Figure 16 presents a heatmap of the $\log$ evidence indices for $100$ randomly sampled genes across the $d$ studies. Here the white shade represents a gene not analyzed by the study while the shade of brown represents an evidence index of 0 which corresponds to a failure to reject the underlying null hypothesis. The heterogeneity across the $d$ studies is evident through the different magnitudes of the evidence indices constructed for each study.
Table 3: Distribution of rejection overlaps across 77 studies.
@@
Since in this example study-specific $p$-values, denoted by $\{p_{ij}\}_{j\in\mathcal{N}_{i},i\in\mathcal{M}}$, are available, one can aggregate the $p$-values pertaining to each hypothesis $i$ and then determine an appropriate threshold for FDR control at level $\alpha$ using the aggregated $p$-values. However, if the underlying model is misspecified the validity of the corresponding $p$-values may be affected. In contrast, $e$-values are relatively more robust to such model misspecification (Wang and Ramdas, 2022) and particularly to dependence between the $p$-values (Vovk and Wang, 2021). So we transform the $p$-values to $e$-values using the following calibrator from Equation (B.1) in Vovk and Wang (2021):
$e_{ij}^{\texttt{P2E}}(\kappa)=\begin{cases}\infty~&\text{if~}p_{ij}=0\\ \dfrac{\kappa(1+\kappa)^{\kappa}}{p_{ij}(-\log p_{ij})^{1+\kappa}}~&\text{if~}p_{ij}\in(0,e^{-\kappa-1}]\\ 0~~&\text{if~}p_{ij}\in(e^{-\kappa-1},1]\end{cases},$
where we choose $\kappa=1$ following the recommendation, and write $e_{ij}^{\texttt{P2E}}:=e_{ij}^{\texttt{P2E}}(1)$. Note that $e_{ij}^{\texttt{P2E}}$ as defined above are bonafide e-values. Therefore, to aggregate $e_{ij}^{\texttt{P2E}}$ we can simply take their average
${e}_{i}^{\texttt{{P2E},agg}}=\dfrac{1}{d}\sum_{j=1}^{d}\Bigl{\{}e_{ij}^{\texttt{P2E}}~\mathbb{I}(i\in\mathcal{M}_{j})+\mathbb{I}(i\notin\mathcal{M}_{j})\Bigr{\}}.$
In this application, we denote the method that applies the e-BH procedure on ${e}_{i}^{\texttt{{P2E},agg}}$ and ${e}_{i}^{\texttt{{P2E},agg*}}$ as P2E and P2E, respectively, and compare them to the inferences obtained from IRT and IRT*.
Ranking and thresholding using IRT and P2E - we aggregate the evidence indices using Equation (4) and threshold the ordered aggregated evidences using the e-BH procedure at $\alpha=0.1$. We recall that this thresholding scheme guarantees valid FDR control under unknown and arbitrary dependence between the aggregated evidences.
@@
The top-left panel of Figure 17 presents a histogram of the log-transformed non-zero aggregated evidence from IRT, while the top-right panel plots the top 25 genes with respect to their aggregated evidence, colored and shape-coded by the number of times the corresponding gene was analyzed across the $d$ studies. Interestingly, the top second and third genes have $n_{i}=2$ and $3$, respectively, suggesting that apart from the number of times a particular null hypothesis is analyzed across the $d$ studies, the magnitude of the study-specific evidence indices also play a key role in the overall ranking. To put this into perspective, the bottom-right panel of Figure 17 presents the top 25 genes with respect to their aggregated evidence from the P2E framework discussed earlier. In stark contrast to IRT, here the top 25 genes have $n_{i}\geq 7$. Furthermore, both the left and right panels of Figure 17 suggest that $e_{i}^{\texttt{P2E}}$ can be substantially larger in magnitude than $e_{i}^{\tt agg}$ particularly when one of the studies rejects the null hypothesis with an astronomically small $p$-value.
@@
Ranking and thresholding using IRT and P2E* - Here we aggregate the evidence indices using the scheme discussed in Section B and threshold the ordered aggregated evidences using the e-BH procedure at $\alpha=0.005$. We note that this thresholding scheme guarantees valid FDR control under (1) exchangeability of the study-specific summary statistics (Definition 2), (2) symmetry of the study-specific decision rule (Definition 3), and (3) independence of the $n_{i}$ summary statistics for each testing problem. In this application the summary statistics are $p$-values which are derived without any side information and so assumption (2) holds. However, verification of assumptions (1) and (3) requires additional information. Nevertheless, Figure 14 reveals that as far as FDR control is concerned, IRT is relatively robust to the violation of the exchangeability assumption (assumption (1)) and for moderate levels of dependence between the $p$-values for each testing problem (assumption (3)), IRT continues to provide valid FDR control.